# Leveraging Large Language Models to Detect Missed Peephole Optimizations

### v1

@@ 4.4. RQ3: Throughput and Cost of Lampo
In summary, these results further demonstrate the feasibility and practical value of Lampo. Although Lampo requires additional computation resources for LLM inference compared to Souper, cost-efficient commercial LLMs (e.g., Gemini2.5 from Google) enables Lampo to achieve high throughput with manageable monetary cost. Moreover, as LLMs continue to become faster and more cost-effective, the overall cost-effectiveness of Lampo will likewise improve.
@@ 4.5. Threats to Validity
The first threat to validity is potential data leakage. When evaluating the effectiveness of Lampo in finding previously reported missed optimization, it is possible that the training set of the deployed model already contains information related to the missed optimizations, which could inflate the results. To mitigate this threat, we selected LLMs with clear knowledge cut-off dates and only included missed optimizations reported after those dates in our benchmarks. Another concern is that the collected previously reported missed optimizations may not be fully representative. To mitigate this, we gathered as many benchmarks as possible and manually inspected each case to ensure validity. The non-determinism of LLMs poses a threat to the internal validity. We mitigate this by evaluating Lampo with multiple LLMs and repeating the controlled experiment in RQ1 for five times.
@@ 5. Related Work
Superoptimization.  A superoptimizer takes a sequence of instructions as input and searches for a more efficient, semantically equivalent or refined program. Massalin (Massalin, 1987) first introduced the concept of superoptimization and developed the earliest superoptimizer based on exhaustive enumeration, which, while pioneering, was both inefficient and unsound. To address scalability limitations, subsequent work proposed alternative search strategies, often leveraging SMT solvers for equivalence verification. Joshi et al. developed Denali (Joshi et al., 2002), a goal-directed superoptimizer that employs expert knowledge to encode equality-preserving transformations, thereby constraining the search space. Bansal and Aiken (Bansal and Aiken, 2006) presented a superoptimizer that enumerates only canonical instruction sequences to reduce redundancy and stores computed optimizations in a database to avoid repeated work. Schkufza et al. introduced STOKE (Schkufza et al., 2013), a superoptimizer for x86-64 binaries that formulates superoptimization as a stochastic search problem and applies Markov Chain Monte Carlo (MCMC) sampling to efficiently explore the vast program space. This approach sacrifices completeness but significantly increases the diversity of candidate programs, thereby improving the quality of the resulting code. Sasnauskas et al. proposed Souper (Sasnauskas et al., 2017), a superoptimizer for LLVM IR that extracts integer-typed functions from a module, traverses backward along dataflow edges from the return instruction, and synthesizes more performant equivalents. Souper is designed to be sound, but its applicability is limited to a purely functional, control-flow-free subset of LLVM IR. Hydra (Mukherjee and Regehr, 2024) addresses the limitations of ungeneralized peephole optimizations produced by superoptimizers by generalizing them using program synthesis. In general, superoptimizers are capable of discovering new optimization patterns, but they are computationally expensive, thus often limited to tiny window size or specific instruction sets. In contrast, Lampo leverages LLMs to explore optimization opportunities, which is less restricted by the code size and instruction sets, and thus can potentially detect more diverse missed optimizations.
@@
Differential Testing for Missed Optimizations.  Differential testing (McKeeman, 1998) compares the outputs of two or more programs on identical inputs to detect discrepancies, which may reveal bugs. Barany (Barany, 2018) applied this technique to compilers, identifying missed optimizations by comparing code generated by different compilers for the same source program. Theodoridis et al. (Theodoridis et al., 2022) extended this approach to dead code elimination, while Liu et al. (Liu et al., 2023) used differential testing to evaluate WebAssembly optimizers by comparing their outputs with those of modern C compilers. Italiano and Cummins (Italiano and Cummins, 2024) further advanced differential testing for code size optimizations, proposing methods such as comparing outputs from seed and mutated code, evaluating different optimization settings, and analyzing outputs from different compiler versions or distinct compilers. Differential testing is effective for discovering missed optimizations, but it is limited to optimizations that have been implemented in other compilers or that are not generalized enough. In contrast, Lampo can potentially discover new optimizations that have not been implemented in any compiler.
@@
LLM-Based Approaches for Optimization.  Recent advances in LLMs have demonstrated their promise in program synthesis, including code generation and optimization tasks. A number of studies have explored the use of these models for optimizations (Garg et al., 2023; Gao et al., 2024; Cummins et al., 2023; Grubisic et al., 2024; Italiano and Cummins, 2024). Some of them focus on source-level code optimization and do not address the problem of finding missed optimizations (Garg et al., 2023; Gao et al., 2024). Cummins et al. (Cummins et al., 2023) investigated the use of LLMs to generate optimized LLVM IR and recommend compiler passes, highlighting their potential but also their limitations in terms of reusability and computational cost. Grubisic et al. (Grubisic et al., 2024) incorporated feedback from optimizer-generated code to enhance optimization quality. The only prior work that directly targets missed optimizations with LLMs is by Italiano and Cummins (Italiano and Cummins, 2024). Their approach employs LLMs to mutate code for differential testing, achieving effective results but restricting the discovery of novel optimization patterns. In contrast, Lampo leverages the code understanding and reasoning capabilities of LLMs to directly detect suboptimal code snippets, which can potentially discovery new unimplemented optimization patterns.
@@ 6. Conclusion
This paper introduced Lampo, a novel framework that bridges the creative capabilities of LLM with the rigorous guarantees of formal verification to tackle the difficult challenge of discovering new peephole optimizations. Our methodology overcomes the limitations of previous approaches by integrating an LLM-based optimizer within a closed-loop system, where verification feedback iteratively refines the LLM’s search for correct and efficient code transformations. Through a comprehensive evaluation on the LLVM compiler, we demonstrated Lampo’s effectiveness in finding a substantial number of previously unreported optimizations. The fact that many of these have been confirmed and fixed in the official LLVM codebase highlights our framework’s practical significance and its ability to outperform state-of-the-art superoptimizers.
Looking ahead, this research opens up exciting new avenues for the use of LLMs in compiler design and beyond. The success of Lampo suggests that similar hybrid methodologies, which combine the generative power of LLMs with rigorous formal methods, could be applied to other complex, pattern-based optimization problems. As the reasoning abilities of LLMs continue to advance, we believe that frameworks like Lampo will become an indispensable tool for continuously improving the performance of modern compilers, paving the way for more efficient and robust software.
@@ Acknowledgements.
This work was partially funded by an unrestricted gift from Google.

### v2

@@ 4.4. RQ3: Throughput and Cost of LPO
In summary, these results further demonstrate the feasibility and practical value of LPO. Although LPO requires additional computation resources for LLM inference compared to Souper, cost-efficient commercial LLMs (e.g., Gemini2.5 from Google) enables LPO to achieve high throughput with manageable monetary cost. Moreover, as LLMs continue to become faster and more cost-effective, the overall cost-effectiveness of LPO will likewise improve.
@@ 4.5. Threats to Validity
The first threat to validity is potential data leakage. When evaluating the effectiveness of LPO in finding previously reported missed optimization, it is possible that the training set of the deployed model already contains information related to the missed optimizations, which could inflate the results. To mitigate this threat, we selected LLMs with clear knowledge cut-off dates and only included missed optimizations reported after those dates in our benchmarks. Another concern is that the collected previously reported missed optimizations may not be fully representative. To mitigate this, we gathered as many benchmarks as possible and manually inspected each case to ensure validity. The non-determinism of LLMs poses a threat to the internal validity. We mitigate this by evaluating LPO with multiple LLMs and repeating the controlled experiment in RQ1 for five times.
5. Discussion
In this section, we discuss the impact of the optimizations discovered by LPO, the key conceptual contribution of the work, and the future work on this research direction.
5.1. Impact of the Discovered Optimizations
In LLVM development practice, each patch implementing a peephole optimization is typically evaluated using LLVM Opt Benchmark (Zheng, 2023), a dataset of LLVM IR files collected from 240 real-world projects written in C, C++ or Rust, and the LLVM compile time tracker (Popov, 2025). Maintainers carefully review each patch and accept it only if: 1 the net effect is positive, as demonstrated in the benchmark; and 2 the compile-time impact is acceptable. In other words, each fixed missed optimization demonstrates tangible benefits in practice anticipated by the community. To date, 13 reported missed optimizations have been fixed with 15 carefully reviewed patch commits, demonstrating that optimizations discovered by LPO have a real-world impact.
Table 5. The Number of impacted IR files and projects and compile time impact for each accepted patch.
ID #IR Files #Projects $\Delta$ in Compile Time
(instruction:u)
128134 54 13 +0.02%
133367 122 18 N/A
142674 251 15 +0.05%
142711 10 1 -0.00%
143211 16 4 N/A
143636 2,476 68 +0.02%
154238 10 4 N/A
157315 6 2 +0.00%
157370 N/A N/A +0.04%
157371 (1) 10 13 N/A
157371 (2) 28 1 +0.02%
157524 N/A N/A -0.03%
163108 (1) 3,055 93 -0.05%
163108 (2) 28 4 -0.01%
166973 759 62 N/A
Prevalence of the Detected Suboptimal Patterns.  We analyzed the evaluation results of the patches that fixed the 13 fixed missed optimization using LLVM Opt Benchmark, recording the number of impacted IR files and projects to estimate the prevalence of the corresponding suboptimal patterns in real-world code. As shown in Table 5, most of these handled suboptimal patterns appear in multiple real-world projects. Notably, Issue 143636 affects 2,476 IR files across 68 projects, and Issue 163108 affects over 3,000 files across 97 projects. These results highlight that LPO can identify impactful missed optimizations.
Impacts on the Compile Time.  For each patch, we also examined the compile-time impact using the LLVM compiler time tracker. Table 5 reports the geometric mean of the compile-time changes across 10 realistic and representative benchmarks (kimwitu++, sqlite3, consumer-typeset, Bullet, tramp3d-v4, mafft, ClamAV, lencod, SPASS, and 7zip) comparing the compiler before and after the patch. The compile-time metric is measured as instruction:u, i.e., the number of machine instructions executed in user space during compilation. A positive change indicates a slowdown. As the table shows, the compile-time impact of each patch is negligible.
Figure 5. Geometric mean of speedup on SPEC CPU2017 Integer for selected patches, as well as for two LLVM versions (f3501d7 on Dec. 1, 2025 vs. 017c75b on Dec. 1, 2024).
Runtime Efficiency on SPEC CPU2017.  To further assess the practical impact of the fixed missed optimizations, we evaluated their runtime effects on the SPEC CPU2017 Integer benchmark suite [^3] We only included C/C++ benchmarks.. We focused on patches that are most likely to affect these benchmarks. [^4] We did not aggregate all patches into a single comparison, as they were committed over time. Cherry-picking them onto a single LLVM version is infeasible due to frequent changes in the related codebase. All experiments were conducted on a system equipped with an AMD Ryzen 9 7950X 16-core processor and 128 GB of RAM. Figure 5 reports the geometric mean speedup. Following the instructions of the benchmark suite, each benchmark was executed three times, and the median runtime was used for comparison (30). The results indicate that none of the evaluated patches yields a significant speedup; all observed performance differences are within 2%, a range where measurement noise can have a substantial impact. We speculate that this limited impact is likely due to the maturity of LLVM. Having been developed and refined for decades, LLVM already incorporates a highly optimized compilation pipeline, making it extremely challenging to achieve substantial performance improvements through a small number of peephole optimizations alone. To further investigate this hypothesis, we compared a recent LLVM version with one released approximately one year earlier; the results are also shown in Figure 5. This comparison similarly reveals no significant runtime performance difference, lending additional support to our speculation.
5.2. The Key Insight and Future Work
The key insight of this work is that, given a reliable and efficient mechanism for rigorously verifying the correctness of optimizations, we can leverage the creativity of LLMs to discover or even directly perform optimizations without being hindered by their inherent unreliability. In this work, we demonstrate the feasibility of this idea by using LLMs to detect missed peephole optimizations in LLVM IR, despite its maturity. Moreover, the fact that LPO is able to identify 62 missed peephole optimizations suggests that state-of-the-art LLMs already possess sufficient code understanding and reasoning capabilities to contribute meaningfully to compiler optimization tasks.
Several promising directions remain for future work. First, although this work focuses on LLVM IR, the underlying idea also applies to other IR (e.g., MLIR) and instruction set architectures (e.g., x86, ARM, and RISC-V). Adapting LPO to these additional targets is therefore an important next step. Second, LPO currently targets only peephole optimizations; future studies could investigate whether LLMs can also help identify missed cases in other classes of optimizations. Third, while LPO is designed to search for missed peephole optimizations, the subsequent steps—generalizing optimization patterns and implementing them in the compiler—are still performed manually by maintainers. Future research could explore how LLMs might also assist with these generalization and implementation phases. Finally, rigorous verification tools are essential to the success of LPO, as they provide strong correctness guarantees for LLM-generated outputs. Accordingly, future work may also focus on developing translation validation tools for additional targets or enhancing existing tools (e.g., Alive2).
@@ 6. Related Work
Superoptimization.  A superoptimizer takes a sequence of instructions as input and searches for a more efficient, semantically equivalent or refined program. Massalin (Massalin, 1987) first introduced the concept of superoptimization and developed the earliest superoptimizer based on exhaustive enumeration, which, while pioneering, was both inefficient and unsound. To address scalability limitations, subsequent work proposed alternative search strategies, often leveraging SMT solvers for equivalence verification. Joshi et al. developed Denali (Joshi et al., 2002), a goal-directed superoptimizer that employs expert knowledge to encode equality-preserving transformations, thereby constraining the search space. Bansal and Aiken (Bansal and Aiken, 2006) presented a superoptimizer that enumerates only canonical instruction sequences to reduce redundancy and stores computed optimizations in a database to avoid repeated work. Schkufza et al. introduced STOKE (Schkufza et al., 2013), a superoptimizer for x86-64 binaries that formulates superoptimization as a stochastic search problem and applies Markov Chain Monte Carlo (MCMC) sampling to efficiently explore the vast program space. This approach sacrifices completeness but significantly increases the diversity of candidate programs, thereby improving the quality of the resulting code. Sasnauskas et al. proposed Souper (Sasnauskas et al., 2017), a superoptimizer for LLVM IR that extracts integer-typed functions from a module, traverses backward along dataflow edges from the return instruction, and synthesizes more performant equivalents. Souper is designed to be sound, but its applicability is limited to a purely functional, control-flow-free subset of LLVM IR. Liu et al. proposed Minotaur (Liu et al., 2024), another superoptimizer for LLVM IR that focuses on optimizations involving integer and floating-point SIMD code. Minotaur supports more operations compared to Souper, but its effectiveness is still constrained by the synthesis-based search strategy. Hydra (Mukherjee and Regehr, 2024) addresses the limitations of ungeneralized peephole optimizations produced by superoptimizers by generalizing them using program synthesis.
Comparison with LPO: Superoptimizers are capable of discovering new optimization patterns, but they are computationally expensive, thus often limited to tiny window size or specific instruction sets. They may also fail to produce results within reasonable time. In contrast, LPO leverages LLMs to explore optimization opportunities, which is less restricted by the code size and instruction sets, and thus can potentially detect more diverse missed optimizations.
@@
Differential Testing for Missed Optimizations.  Differential testing (McKeeman, 1998) compares the outputs of two or more programs on identical inputs to detect discrepancies, which may reveal bugs. Barany (Barany, 2018) applied this technique to compilers, identifying missed optimizations by comparing code generated by different compilers for the same source program. Theodoridis et al. (Theodoridis et al., 2022) extended this approach to dead code elimination, while Liu et al. (Liu et al., 2023) used differential testing to evaluate WebAssembly optimizers by comparing their outputs with those of modern C compilers. Italiano and Cummins (Italiano and Cummins, 2024) further advanced differential testing for code size optimizations, proposing methods such as comparing outputs from seed and mutated code, evaluating different optimization settings, and analyzing outputs from different compiler versions or distinct compilers.
Comparison with LPO: Differential testing is effective for discovering missed optimizations, but it is limited to optimizations that have been implemented in other compilers or that are not generalized enough. In contrast, LPO can potentially discover new optimizations that have not been implemented in any compiler.
@@
LLM-Based Approaches for Optimization.  Recent advances in LLMs have demonstrated their promise in program synthesis, including code generation and optimization tasks. A number of studies have explored the use of these models for optimizations (Garg et al., 2023; Gao et al., 2024; Cummins et al., 2023; Grubisic et al., 2024; Italiano and Cummins, 2024). Some of them focus on source-level code optimization and do not address the problem of finding missed optimizations (Garg et al., 2023; Gao et al., 2024). Cummins et al. (Cummins et al., 2023) investigated the use of LLMs to generate optimized LLVM IR and recommend compiler passes, highlighting their potential but also their limitations in terms of reusability and computational cost. Grubisic et al. (Grubisic et al., 2024) incorporated feedback from optimizer-generated code to enhance optimization quality.
Comparison with LPO: The only prior work that aims to detect missed optimizations with LLMs is by Italiano and Cummins (Italiano and Cummins, 2024). Their approach employs LLMs to mutate code for differential testing, achieving effective results but restricting the discovery of novel optimization patterns. In contrast, LPO leverages the code understanding and reasoning capabilities of LLMs to directly detect suboptimal code snippets, which can potentially discover new unimplemented optimization patterns.
@@ 7. Conclusion
This paper introduced LPO, a novel framework that bridges the creative capabilities of LLM with the rigorous guarantees of formal verification to tackle the difficult challenge of discovering new peephole optimizations. Our methodology overcomes the limitations of previous approaches by integrating an LLM-based optimizer within a closed-loop system, where verification feedback iteratively refines the LLM’s search for correct and efficient code transformations. Through a comprehensive evaluation on the LLVM compiler, we demonstrated LPO’s effectiveness in finding a substantial number of previously unreported optimizations. The fact that many of these have been confirmed and fixed in the official LLVM codebase highlights our framework’s practical significance and its ability to outperform state-of-the-art superoptimizers.
Looking ahead, this research opens up exciting new avenues for the use of LLMs in compiler design and beyond. The success of LPO suggests that similar hybrid methodologies, which combine the generative power of LLMs with rigorous formal methods, could be applied to other complex, pattern-based optimization problems. As the reasoning abilities of LLMs continue to advance, we believe that frameworks like LPO will become an indispensable tool for continuously improving the performance of modern compilers, paving the way for more efficient and robust software.
@@ Acknowledgements.
We thank all the anonymous reviewers of ASPLOS’26 for their helpful feedback and constructive comments. We also thank the LLVM maintainers and contributors, especially Yingwei Zheng, for their efforts in addressing the reported missed optimizations and for their valuable feedback. This research is partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant, CFI-JELF Project #40736, and an unrestricted gift from Google.